OpenStudy (anonymous):

Which of the following is equivalent to five and four-ninths divided by seven? seven-ninths one and two-sevenths five-sevenths five and four sixty-thirds

OpenStudy (anonymous):

@Chineseboy15

OpenStudy (anonymous):

seven ninths

OpenStudy (anonymous):

five and four ninths = 49/9 49/9 divided by 7 = 49/9 * 1/7 = 49/63 = 7/9